1) The probability that a graduating student will get an offer for job A is 0.35 and the probability that this student will get an offer for job B is 0.20. If the probability of this student getting both job offers is 0.05, what is the probability that this student will get at least one of the two job offers?
Answer by ikleyn(52847)

.

    P(at least one) = P(A or B) = P(A) + P(B) - P(A and B) = 0.35 + 0.2 - 0.05.